OpenStudy (anonymous):

The radius of a circle is rt=t^3+t^2+1 inches at time t minutes for 0≺t≺10. The radius of the circle is changing at the rate _______when t=.5 minutes. The circumference of the circle is changing at the rate ________when t=.5 minutes. The area of the circle is changing at the rate_______ when t=.5 minutes.

OpenStudy (dumbcow):

im assuming radius is r. r = t^3 +t^2 +1 Find dr/dt by taking derivative dr/dt = 3t^2 +2t circumference is 2*pi*r c = 2*pi(t^3 +t^2 +1) Take derivative dc/dt = 2*pi(3t^2 +2t) area is pi*r^2 A = pi [t^3 +t^2 +1] ^2 dA/dt = 2pi(t^3 +t^2 +1)(3t^2+ 2t)